In a survey of 7000 women, 4431 say they change their polish once a week. Construct a 99% confidence interval for the population proportion of women who change their polish once a week.
Answer by stanbon(75887)

In a survey of 7000 women, 4431 say they change their polish once a week. Construct a 99% confidence interval for the population proportion of women who change their polish once a week.
-------
p-hat = 4431/7000 = 0.633
----
E = 2.5758*sqrt[0.633*0.367/7000] = 0.0148
===============
99% CI: 0.633- 0.0148 < p < 0.633+ 0.0148
----
99% CI: 0.6182 < p < 0.6478
================
Cheers,
Stan H.
================